OpenStudy (anonymous):

is (2,3) a soluition of the system 3x+4y=18 2x-y=1

OpenStudy (anonymous):

Yes, it is ;)

OpenStudy (anonymous):

Sobstitute x with 2 in first(or second) equation and y with 3 in the same equation, if it's an equivalence then it's true ;)

OpenStudy (anonymous):

yes, (2) 3x+4y=18 (3) 2x-y=1 6x+8y=36 -6x-3y=3 =8y=36 +-3y=3 y=3 2x-(3)=1 2x=4 x=2

OpenStudy (anonymous):

THANK YOU BOTH I JUST HAVE TO PLUF IN THE Y AND X =)